OpenStudy (anonymous):

Quadrilateral QRST is a square. If the measure of angle RQS is 5x – 15. Find the value of x.

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

Every angle in a square is 90 degree. So, 5x-15=90 Solve for x.

OpenStudy (anonymous):

igot x=21 ... is that right?

OpenStudy (aravindg):

Yep :)

OpenStudy (anonymous):

are u sure ?

OpenStudy (aravindg):

RQS=1/2(RQT)=1/2(90)=45

OpenStudy (anonymous):

x=45?